# Overcoming Saturation in Density Ratio Estimation by Iterated Regularization

### v1

@@ 1 Introduction
Given two i.i.d. samples ${\bf x}=\{x_{i}\}_{i=1}^{m}$ and ${\bf x}^{\prime}=\{x_{i}^{\prime}\}_{i=1}^{n}$, drawn from two probability measures $P$ and $Q$, respectively, the problem of density ratio estimation is to approximate $\beta:=\frac{\mathop{}\!\mathrm{d}P}{\mathop{}\!\mathrm{d}Q}$. This problem arises in anomaly detection (Smola et al., 2009; Hido et al., 2011), two-sample testing (Keziou and Leoni-Aubin, 2005; Kanamori et al., 2011), divergence estimation (Nguyen et al., 2007, 2010b), unsupervised domain adaptation (Shimodaira, 2000; Dinu et al., 2023), generative modeling (Mohamed and Lakshminarayanan, 2016), conditional density estimation (Schuster et al., 2020), and PU learning (Kato et al., 2019).
A large class of methods for density ratio estimation relies on the estimation of minimizers of the form Sugiyama et al. (2012b, a)
@@
  • The logistic regression approach (LR) employed in (Bickel et al., 2009, Section 7) can be obtained using $F(h)=\int_{\mathcal{X}}h(x)\log(h(x))-(1+h(x))\log(1+h(x))\mathop{}\!\mathrm{d}Q(x)$ and $g(f)=e^{f}$.
  • The exponential approach (Exp) in Menon and Ong (2016) is realized by $F(h)=\int_{\mathcal{X}}h(x)^{-3/2}\mathop{}\!\mathrm{d}Q(x)$, $g(f)=e^{2f}$.
  • The square loss approach (SQ) in Menon and Ong (2016) originates from $F(h)=\int_{X}1/(2h(x)+2)\mathop{}\!\mathrm{d}Q(x)$, $g(f)=\frac{-1+2f}{2-2f}$.
However, state-of-the-art error guarantees as Kanamori et al. (2012b); Que and Belkin (2013); Gizewski et al. (2022); Zellinger et al. (2023) for the methods above are sub-optimal for problems of high regularity, i.e., they don’t achieve optimal convergence rates when increasing the number of samples, see Theorem 1 and Figure 1. This issue is called saturation in learning theory (Bauer et al., 2007; Beugnot et al., 2021; Li et al., 2022) and inverse problems (Engl et al., 1996), and its negative effect in supervised learning has been recently highlighted (Beugnot et al., 2021).
In this work, we approach the problem of saturation in density ratio estimation by iterated regularization (cf. King and Chillingworth (1979); Beugnot et al. (2021)) of the form
@@
with $f^{\lambda,0}=0$. We prove that iterative methods, according to Eq. (2), do not suffer from saturation. For $t=1$, Eq. (2) includes, e.g., approaches in the example above. As a consequence, we are able to improve the empirical performance of these approaches on benchmark experiments, using a higher iteration number $t>1$ ($t\leq 10$ in our experiments).
@@
  • We provide the (to the best of our knowledge first) non-saturating error bounds for density ratio estimation, which are optimal for the square loss approach Menon and Ong (2016) in the sense of Caponnetto and De Vito (2007).
@@
Figure 1: Saturation issue of classical methods Eq. (1) versus our novel iteratively regularized approach Eq. (2). Left: Error rates proven in Theorem 1 for classical methods (red) and ours (blue). Right: Error for classical KuLSIF Kanamori et al. (2009) method with $F(h)=\int(h(x)-1)^{2}/2q(x)\mathop{}\!\mathrm{d}x$ in Eq. (1) (red), and our iteratively regularized approach (blue) applied to Gaussian mixture $P$ and Gaussian distribution $Q$; smaller number of components allows higher regularity index $r$. The residual differences (grey) between the methods increase with higher regularity.
@@ 2 Related Work
It was first observed in Sugiyama et al. (2012b) that many loss functions for density ratio estimation can be modeled by a Bregman divergence $B_{F}(\beta,g(f))$ as used in Eq.(1), see also Menon and Ong (2016); Kato and Teshima (2021) and references therein.
There are some theoretical results on the large class of algorithms following variations of Eq. (1), see Kanamori et al. (2012b); Kato and Teshima (2021); Gizewski et al. (2022); Zellinger et al. (2023). However, to the best of our knowledge, none of these works achieves the optimal error rate discovered in Caponnetto and De Vito (2007) for supervised learning (for $r>\frac{1}{2}$ in Assumption 1). We are also not aware of any approach provably overcoming the saturation problem of Tikhonov regularization $\left\lVert f\right\rVert^{2}$ in Eq. (1), except Nguyen et al. (2023) who study iterated Lavrentiev regularization only for the special case of KuLSIF. They include only a small numerical visualization and their pointwise error measure is different from our $B_{F}$. Our approach overcomes saturation and provably achieves the optimal rate of square loss.
Our proof extends fast error rates from kernel regression (Beugnot et al., 2021) to density ratio estimation, using the duality between divergence estimation (Reid and Williamson, 2011) and binary classification (Menon and Ong, 2016), and the self-concordance property of associated loss functions (Bach, 2010; Marteau-Ferey et al., 2019b).
@@
The problem of saturation has been studied in inverse problems (Engl et al., 1996), in a general form in learning theory (Bauer et al., 2007), and recently for general (self-concordant) convex loss functions in Beugnot et al. (2021). Our work can be regarded as an extension of these investigations from supervised learning to density ratio estimation.
@@
It is important to note that iterated Tikhonov regularization is intrinsically related to the proximal point method (Rockafellar, 1976). We also mention the related line of works Rhodes et al. (2020); Choi et al. (2021) for problems where $P$ and $Q$ are very dis-similar, which can be combined with our methods.
@@
In our empirical studies, we rely on the benchmark experiments for density ratio estimation of Kanamori et al. (2012b) and the large scale-domain adaptation framework Ragab et al. (2023) with importance weighting based parameter choice as introduced in Dinu et al. (2023).
@@ 3 Notation
Let us denote by $\mathcal{X}$ a (compact) input space and by $\mathcal{Y}:=\{-1,1\}$ the binary label space and assume that an input $x$ is drawn with probability [^1] See the supplementary material of Menon and Ong (2016) for probabilities $\pi\neq\frac{1}{2}$. $\frac{1}{2}$ from $P$ and $Q$. That is, there is a probability measure $\rho$ on $\mathcal{X}\times\mathcal{Y}$, with conditional measures $\rho(x|y=1):=P(x),\rho(x|y=-1):=Q(x)$ and marginal measure $\rho_{\mathcal{Y}}$ defined as Bernoulli measure assigning probability $\frac{1}{2}$ to both events $y=1,y=-1$, see Reid and Williamson (2010). That is, the sample ${\bf z}:=(x_{i},1)_{i=1}^{m}\cup(x_{i}^{\prime},-1)_{i=1}^{n}$ can be regarded as an i.i.d. sample of an $\mathcal{X}\times\mathcal{Y}$-valued random variable $Z$ with measure $\rho$.
@@ 4.1 Algorithmic Realization
To solve Eq. (4), we follow the representer theorem, make the model ansatz $f(\cdot)=\sum_{i=1}^{m+n}\alpha_{i}k(x_{i},\cdot)$ and apply the conjugate gradient (CG) method for minimization, except for KuLSIF in Example 2, for which an explicit solution is available, see Appendix C. Originally, the CG method was developed to solve linear equations with a positive semi-definite matrix. There, the goal is to compute a sequence of conjugate search directions, where two vectors are considered conjugate if their dot product with respect to the given matrix is zero. These search directions arise as gradients of an associated quadratic optimization problem, see e.g. Golub and Van Loan (2013, Section 11.3).
The CG-method can be generalized to deal with nonlinear problems like Eq. (4), which leads to a minimization problem of a strongly convex smooth function in each iteration step. Therefore, several CG approaches can be applied, see, e.g., Hager and Zhang (2006). The method which is implemented in Python Scipy (Virtanen et al., 2020) and which we are using, is based on an additional Polak-Ribiere line search scheme and it is guaranteed to converge globally in our case (Hager and Zhang, 2006; Cohen, 1972). In the upcoming section we will also briefly discuss how this algorithm choice affects our theoretical guarantees.
@@ 5 Learning Theoretic Analysis
In learning theory, the regularity of learning problems is traditionally encoded by source conditions, see, e.g., Bauer et al. (2007); Caponnetto and De Vito (2007); Marteau-Ferey et al. (2019b).
@@
To interpret Assumption 1, let us assume $\ell_{z}(f):=(y-f(x))^{2}$ is the square loss. Then Assumption 1 recovers the polynomial source condition used in Caponnetto and De Vito (2007). Starting from $r=0$, the regularity of $f_{\mathcal{H}}$ increases with increasing $r$, essentially reducing the number of eigenvectors of the covariance operator $T$ needed to well approximate $f_{\mathcal{H}}$. However, it is known that optimal learning rates are only possible under a further condition on the capacity of the space $\mathcal{H}$ (cf. Marteau-Ferey et al. (2019b)).
@@
The degrees of freedom term $\mathrm{df}_{\lambda}$ is also called Fisher information (Van der Vaart, 2000), appears in spline smoothing (Wahba, 1990), and reduces to the effective dimension (Caponnetto et al., 2005) for the square loss. In the latter case, a bigger $\alpha$ implies that fewer eigenvectors are needed to approximate elements from the image space of $T$ with a given capacity, cf. Blanchard and Mücke (2018).
Of course, it is not possible to achieve fast convergence with any loss function. We therefore focus on the large class of convex pseudo self-concordant losses, see Bach (2010); Ostrovskii and Bach (2021).
@@
For $r\leq\frac{1}{2},t=1$, Theorem 1 recovers the state-of-the-art error rate of Zellinger et al. (2023). In addition, it extends the state of the art for current approaches for $r>\frac{1}{2},t=1$. In case of square loss, it is optimal (cf. Caponnetto and De Vito (2007); Blanchard and Mücke (2018)). For $r>\frac{1}{2},t>1$, Theorem 1 provides the first non-saturating error rate for density ratio estimation.
@@
The parameter choice in Eq. (7) balances a bias and a variance term in a tight error bound (Marteau-Ferey et al., 2019b) and is thus rate optimal, see, e.g., Lu and Pereverzev (2013, Section 4.3). Eq. (7) is a priori and requires the knowledge of the regularity index $r$. This issue can be resolved using an a posteriori parameter choice rule, e.g., De Vito et al. (2010); Zellinger et al. (2023).
@@
For nonlinear CG algorithms and strictly convex smooth objectives, $\log(\varepsilon^{-1})$ gradient and function evaluations are required to achieve an error tolerance $\varepsilon$, see e.g. Neumaier et al. (2022); Chan-Renous-Legoubin and Royer (2022). Combining these findings with  arguments from Beugnot et al. (2021, Section 3.2), we see that $\varepsilon$ influences Theorem 1 in an additive way, i.e. with probability at least $1-\delta$,
@@ 6 Empirical Evaluations
We investigate the performance of iterated density ratio estimation regarding three aspects below. Please find further details in Appendix E.
@@ 6.1 Datasets
To illustrate Theorem 1, we adapt an example of Beugnot et al. (2021, Section 4) to density ratio estimation. In this example, the regularity index $r$ and the capacity parameter $\alpha$ are known by design. Let therefore $\mathcal{X}=[0,1]$ and the kernel of $\mathcal{H}$ given by $k(x,y)=h_{\alpha}(x,y)=1+\sum_{l\in\mathbb{Z}\setminus 0}\frac{e^{2il\pi(x-y)}}{|l|^{\alpha}}$, for which an explicit formula is known for even $\alpha$ (Wahba, 1990, Page 22), which is essentially determined by Bernoulli polynomials (Olver et al., 2010, Section 24.2). If we fix $\rho_{\mathcal{X}}$ as the uniform distribution, $\ell$ the logistic loss and $f_{\mathcal{H}}(x)=h_{(r+\frac{1}{2})\alpha+\frac{1}{2}}(0,x)$, then $f_{\mathcal{H}}\in\mathcal{H}$ and Assumptions 1 and 2 are satisfied (Rudi and Rosasco, 2017). Moreover, it is shown in Ciliberto et al. (2020, Lemma A1) that
@@
resulting in the density ratio $\beta(x)=\frac{1-\pi}{\pi}\frac{\rho(y=1|x)}{1-\rho(y=1|x)}$ and a different $B_{F}$, see Menon and Ong (2016, Section 4.2). We investigate the sample convergence in the distribution-independent error $\left\lVert\beta-g(f_{\bf z}^{\lambda,t})\right\rVert_{L^{1}([0,1])}$.
@@
The DomainNet-2019 dataset (Peng et al., 2019) consists of six different image domains (Quickdraw: Q, Real: R, Clipart: C, Sketch: S, Infograph: I, and Painting: P). We follow Zellinger et al. (2021) and rely on the reduced version of DomainNet-2019, referred to as MiniDomainNet, which reduces the number of classes to the top five largest representatives in the training set across all six domains. To further improve computation time, we rely on a ImageNet (Krizhevsky et al., 2012) pre-trained ResNet-18 (He et al., 2016) backbone. Therefore, we assume that the backbone has learned lower-level filters suitable for the "real" image category, and we only need to adapt to the remaining five domains (Clipart, Sketch,…). This results in five domain adaptation tasks.
@@
The Heterogeneity Human Activity Recognition (Stisen et al., 2015) dataset investigate sensor-, device- and workload-specific heterogeneities using $36$ smartphones and smartwatches, consisting of $13$ different device models from four manufacturers. In total, we used all five domain adaptation scenarios from Dinu et al. (2023, Tables 25 and 26).

### v2

@@ 1 Introduction
Given two i.i.d. samples ${\bf x}=\{x_{i}\}_{i=1}^{m}$ and ${\bf x}^{\prime}=\{x_{i}^{\prime}\}_{i=1}^{n}$, drawn from two probability measures $P$ and $Q$, respectively, the problem of density ratio estimation is to approximate $\beta:=\frac{\mathop{}\!\mathrm{d}P}{\mathop{}\!\mathrm{d}Q}$. This problem arises in anomaly detection (Smola et al., 2009; Hido et al., 2011), two-sample testing (Keziou & Leoni-Aubin, 2005; Kanamori et al., 2011), divergence estimation (Nguyen et al., 2007, 2010b), unsupervised domain adaptation (Shimodaira, 2000; Dinu et al., 2023), generative modeling (Mohamed & Lakshminarayanan, 2016), conditional density estimation (Schuster et al., 2020), and PU learning (Kato et al., 2019).
A large class of methods for density ratio estimation relies on the estimation of minimizers of the form Sugiyama et al. (2012b, a); Zellinger et al. (2023)
@@
In the following, we exemplify four methods in this class. We refer to Sugiyama et al. (2012a, Section 7) and Menon & Ong (2016); Kato & Teshima (2021) for further examples.
@@
  • The logistic regression approach (LR) employed in Bickel et al. (2009, Section 7) can be obtained using $F(h)=\int_{\mathcal{X}}h(x)\log(h(x))-(1+h(x))\log(1+h(x))\mathop{}\!\mathrm{d}Q(x)$ and $g(f)=e^{f}$.
  • The exponential approach (Exp) in Menon & Ong (2016) is realized by $F(h)=\int_{\mathcal{X}}h(x)^{-3/2}\mathop{}\!\mathrm{d}Q(x)$, $g(f)=e^{2f}$.
  • The square loss approach (SQ) in Menon & Ong (2016) originates from $F(h)=\int_{X}1/(2h(x)+2)\mathop{}\!\mathrm{d}Q(x)$, $g(f)=\frac{-1+2f}{2-2f}$.
However, state-of-the-art error guarantees as Kanamori et al. (2012b); Que & Belkin (2013); Gizewski et al. (2022); Zellinger et al. (2023) for the methods above are sub-optimal for problems of high regularity, i.e., they don’t achieve optimal convergence rates when increasing the number of samples, see Theorem 1 and Figure 1. This issue is called saturation in inverse problems (Engl et al., 1996b) and learning theory (Bauer et al., 2007; Li et al., 2022), and its negative effect in supervised learning has been recently highlighted (Beugnot et al., 2021).
Inspired by the Bregman method (Bregman, 1967) and Bregmanization (Osher et al., 2005; Resmerita & Scherzer, 2006), we approach saturation of Eq. (1) by iterated regularization (King & Chillingworth, 1979) of the form
@@
with $f^{\lambda,0}=0$. We prove that iterative density ratio estimation methods, according to Eq. (2), do not suffer from saturation. For $t=1$, Eq. (2) includes, e.g., approaches in the example above. As a consequence, we are able to improve the empirical performance of these approaches on benchmark experiments, using a higher iteration number $t>1$ ($t\leq 10$ in our experiments).
@@
  • We provide the (to the best of our knowledge first) non-saturating error bounds for density ratio estimation, which are optimal for the square loss approach (Menon & Ong, 2016) in the sense of Caponnetto & De Vito (2007).
@@
Figure 1: Saturation issue of classical methods Eq. (1) versus our novel iteratively regularized approach Eq. (2). Left: Error rates proven in Theorem 1 for classical methods (red) and ours (blue). Right: Error for classical KuLSIF (Kanamori et al., 2009) method with $F(h)=\int(h(x)-1)^{2}/2q(x)\mathop{}\!\mathrm{d}x$ in Eq. (1) (red), and our iteratively regularized approach (blue) applied to Gaussian mixture $P$ and Gaussian distribution $Q$; smaller number of components allows higher regularity index $r$. The residual differences (grey) between the methods increase with higher regularity.
@@ 2 Related Work
It was first observed in Sugiyama et al. (2012b) that many loss functions for density ratio estimation can be modeled by a Bregman divergence $B_{F}(\beta,g(f))$ as used in Eq. (1), see also Menon & Ong (2016); Kato & Teshima (2021) and references therein.
@@
The study of saturation issues in regularization dates back to inverse problems literature (Neubauer, 1997; Engl et al., 1996b) and has been further investigated in learning theory (Bauer et al., 2007; Caponnetto & De Vito, 2007; Li et al., 2022; Lin et al., 2020); which includes recent studies for general (self-concordant) convex loss functions (Beugnot et al., 2021) and vector-valued output spaces (Meunier et al., 2024). Our work can be regarded as an extension of many of these investigations from supervised learning to density ratio estimation.
There are some theoretical results on the large class of algorithms following variations of Eq. (1), see Kanamori et al. (2012b); Kato & Teshima (2021); Gizewski et al. (2022); Zellinger et al. (2023). However, to the best of our knowledge, none of these works achieves the optimal error rate discovered in Caponnetto & De Vito (2007) for supervised learning (for $r>\frac{1}{2}$ in Assumption 1). We are also not aware of any approach provably overcoming the saturation problem of Tikhonov regularization $\left\lVert f\right\rVert^{2}$ in Eq. (1), except Nguyen et al. (2023) who study iterated Lavrentiev regularization only for the special case of KuLSIF. They include only a small numerical visualization and their pointwise error measure is different from our $B_{F}$. Our approach overcomes saturation and provably achieves the optimal rate of square loss.
As noted above, Eq. (2) is inspired by the Bregman method (Bregman, 1967) and Bregmanization (Osher et al., 2005; Resmerita & Scherzer, 2006) as studied in Inverse Problems. It is demonstrated in Resmerita & Scherzer (2006), that regularization by Bregman distance can be considered as an instance of enhancing techniques, commonly used in image processing. We refer to the book (Scherzer et al., 2009) for in-depth convergence studies.
Eq. (4) identifies the proposed methods as iterated Tikhonov regularizations, which have been studied in Inverse Problems literature, e.g., for linear problems (Hanke & Groetsch, 1998), for non-linear problems (Scherzer, 1993), in the continuous case (Scherzer & Groetsch, 2001), for non-quadratic penalties (Tadmor et al., 2004; Kindermann et al., 2023) and for non-linear operators combined with non-linear penalty (Modin et al., 2019). All these studies can serve as a basis for analysing Eq. (2).
Our analysis of the sampling behaviour extends fast error rates from kernel regression (Beugnot et al., 2021) (which relies on Engl et al. (1996a); Bauer et al. (2007)) to density ratio estimation, using the duality between divergence estimation (Reid & Williamson, 2011) and binary classification (Menon & Ong, 2016), and the self-concordance property of associated loss functions (Bach, 2010; Marteau-Ferey et al., 2019b).
@@
It is important to note that iterated Tikhonov regularization is intrinsically related to the proximal point method (Rockafellar, 1976).
The focus of our work is to improve density ratio estimation for methods with theoretical guarantees such as kernel methods and infinitely wide neural networks, see Jacot et al. (2018). For general deep learning, our theory does not apply, as additional regularizations appear (Gidel et al., 2019). However, some recent works identify scientific deep density ratio estimation problems. For example Rhodes et al. (2020) and Kato & Teshima (2021); Kiryo et al. (2017) (PU-learning) and Srivastava et al. (2023) discuss problems (e.g., density-chasm) appearing for very different densities; note that saturation also happens for very similar densities (see Figure 2). Notably, Choi et al. (2021) propose an approach by training normalizing flows to obtain closer and simpler densities. Another important recent line of research is density ratio estimation for correcting deep generative diffusion models (Kim et al., 2023). Notably, in Kim et al. (2024), the Bregman divergence approach in Eq. (1) is extended to time-dependent correction and they identify further interesting Bregman divergences.
@@
In our empirical studies, we rely on the benchmark experiments for density ratio estimation of Kanamori et al. (2012b) and the large scale-domain adaptation framework of Ragab et al. (2023) with importance weighting based parameter choice as introduced in Dinu et al. (2023).
@@ 3 Notation
Let us denote by $\mathcal{X}$ a (compact) input space and by $\mathcal{Y}:=\{-1,1\}$ the binary label space and assume that an input $x$ is drawn with probability [^1] See the supplementary material of (Menon & Ong, 2016) for probabilities $\pi\neq\frac{1}{2}$. $\frac{1}{2}$ from $P$ and $Q$. That is, there is a probability measure $\rho$ on $\mathcal{X}\times\mathcal{Y}$, with conditional measures $\rho(x|y=1):=P(x),\rho(x|y=-1):=Q(x)$ and marginal measure $\rho_{\mathcal{Y}}$ defined as Bernoulli measure assigning probability $\frac{1}{2}$ to both events $y=1,y=-1$, see Reid & Williamson (2010). That is, the sample ${\bf z}:=(x_{i},1)_{i=1}^{m}\cup(x_{i}^{\prime},-1)_{i=1}^{n}$ can be regarded as an i.i.d. sample of an $\mathcal{X}\times\mathcal{Y}$-valued random variable $Z$ with measure $\rho$.
@@ 4.1 Algorithmic Realization
To solve Eq. (4), we follow the representer theorem (Wahba, 1990; Schölkopf et al., 2001) , make the model ansatz $f(\cdot)=\sum_{i=1}^{m+n}\alpha_{i}k(x_{i},\cdot)$ and apply the conjugate gradient (CG) method for minimization, except for KuLSIF in Example 2, for which an explicit solution is available, see Appendix C. Originally, the CG method was developed to solve linear equations with a positive semi-definite matrix. There, the goal is to compute a sequence of conjugate search directions, where two vectors are considered conjugate if their dot product with respect to the given matrix is zero. These search directions arise as gradients of an associated quadratic optimization problem, see e.g. Golub & Van Loan (2013, Section 11.3).
The CG-method can be generalized to deal with nonlinear problems like Eq. (4), which leads to a minimization problem of a strongly convex smooth function in each iteration step. Therefore, several CG approaches can be applied, see, e.g., Hager & Zhang (2006). The method which is implemented in Python Scipy (Virtanen et al., 2020) and which we are using, is based on an additional Polak-Ribiere line search scheme and it is guaranteed to converge globally in our case (Hager & Zhang, 2006; Cohen, 1972). In the upcoming section we will also briefly discuss how this algorithm choice affects our theoretical guarantees.
@@ 5 Learning Theoretic Analysis
In learning theory, the regularity of learning problems is traditionally encoded by source conditions, see, e.g., Bauer et al. (2007); Caponnetto & De Vito (2007); Marteau-Ferey et al. (2019b).
@@
To interpret Assumption 1, let us assume $\ell_{z}(f):=(y-f(x))^{2}$ is the square loss. Then Assumption 1 recovers the polynomial source condition used in Caponnetto & De Vito (2007). Starting from $r=0$, the regularity of $f_{\mathcal{H}}$ increases with increasing $r$, essentially reducing the number of eigenvectors of the covariance operator $T$ needed to well approximate $f_{\mathcal{H}}$. However, it is known that optimal learning rates are only possible under a further condition on the capacity of the space $\mathcal{H}$ (Marteau-Ferey et al., 2019b).
@@
The degrees of freedom term $\mathrm{df}_{\lambda}$ is also called Fisher information (Van der Vaart, 2000), appears in spline smoothing (Wahba, 1990), and reduces to the effective dimension (Caponnetto et al., 2005) for the square loss. In the latter case, a bigger $\alpha$ implies that fewer eigenvectors are needed to approximate elements from the image space of $T$ with a given capacity (Blanchard & Mücke, 2018).
Of course, it is not possible to achieve fast convergence with any loss function. We therefore focus on the large class of convex pseudo self-concordant losses, see Bach (2010); Ostrovskii & Bach (2021).
@@
For $r\leq\frac{1}{2},t=1$, Theorem 1 recovers the state-of-the-art error rate of Zellinger et al. (2023). In addition, it extends the state of the art for current approaches for $r>\frac{1}{2},t=1$. In case of square loss, it is optimal (cf. Caponnetto & De Vito (2007); Blanchard & Mücke (2018)). For $r>\frac{1}{2},t>1$, Theorem 1 provides the first non-saturating error rate for density ratio estimation.
@@
The parameter choice in Eq. (7) balances a bias and a variance term in a tight error bound (Marteau-Ferey et al., 2019b) and is thus rate optimal, see, e.g., Lu & Pereverzev (2013, Section 4.3). Eq. (7) is a priori and requires the knowledge of the regularity index $r$. This issue can be resolved using an a posteriori parameter choice rule, e.g., De Vito et al. (2010); Zellinger et al. (2023).
@@
For nonlinear CG algorithms and strictly convex smooth objectives, $\log(\varepsilon^{-1})$ gradient and function evaluations are required to achieve an error tolerance $\varepsilon$, see e.g. Neumaier et al. (2022); Chan-Renous-Legoubin & Royer (2022). Combining these findings with  arguments from Beugnot et al. (2021, Section 3.2), we see that $\varepsilon$ influences Theorem 1 in an additive way, i.e. with probability at least $1-\delta$,
@@ 6 Empirical Evaluations
We investigate the performance of iterated density ratio estimation regarding three aspects below. We include additional experiments conducted during the rebuttal period that helped to improve our work. More precisely, we compare to a SOTA domain adaptation method (Dinu et al., 2023), combine our approach with telescoping from Rhodes et al. (2020), and extend our approach to Deep Learning methods. Please find further details in Appendix E.
@@ 6.1 Datasets
To illustrate Theorem 1, we adapt an example of Beugnot et al. (2021, Section 4) to density ratio estimation. In this example, the regularity index $r$ and the capacity parameter $\alpha$ are known by design. Let therefore $\mathcal{X}=[0,1]$ and the kernel of $\mathcal{H}$ given by $k(x,y)=h_{\alpha}(x,y)=1+\sum_{l\in\mathbb{Z}\setminus 0}\frac{e^{2il\pi(x-y)}}{|l|^{\alpha}}$, for which an explicit formula is known for even $\alpha$ (Wahba, 1990, Page 22), which is essentially determined by Bernoulli polynomials (Olver et al., 2010, Section 24.2). If we fix $\rho_{\mathcal{X}}$ as the uniform distribution, $\ell$ the logistic loss and $f_{\mathcal{H}}(x)=h_{(r+\frac{1}{2})\alpha+\frac{1}{2}}(0,x)$, then $f_{\mathcal{H}}\in\mathcal{H}$ and Assumptions 1 and 2 are satisfied (Rudi & Rosasco, 2017). Moreover, it is shown in Ciliberto et al. (2020, Lemma A1) that
@@
resulting in the density ratio $\beta(x)=\frac{1-\pi}{\pi}\frac{\rho(y=1|x)}{1-\rho(y=1|x)}$ and a different $B_{F}$, see Menon & Ong (2016, Section 4.2). We investigate the sample convergence in the distribution-independent error $\left\lVert\beta-g(f_{\bf z}^{\lambda,t})\right\rVert_{L^{1}([0,1])}$.
@@
The DomainNet-2019 dataset (Peng et al., 2019) consists of six different image domains (Quickdraw: Q, Real: R, Clipart: C, Sketch: S, Infograph: I, and Painting: P). We follow Zellinger et al. (2021) and rely on the reduced version of DomainNet-2019, referred to as MiniDomainNet, which reduces the number of classes to the top five largest representatives in the training set across all six domains. To further improve computation time, we rely on a ImageNet (Krizhevsky et al., 2012) pre-trained ResNet-18 (He et al., 2016) backbone. Therefore, we assume that the backbone has learned lower-level filters suitable for the ”real” image category, and we only need to adapt to the remaining five domains (Clipart, Sketch,…). This results in five domain adaptation tasks.
@@
The Heterogeneity Human Activity Recognition (Stisen et al., 2015) dataset investigate sensor-, device- and workload-specific heterogeneities using $36$ smartphones and smartwatches, consisting of $13$ different device models from four manufacturers. In total, we used all fiv domain adaptation scenarios from Dinu et al. (2023, Tables 25 and 26).